OpenStudy (anonymous):

Write the equation of a hyperbola with vertices (3, -1) and (3, -9) and co-vertices (-6. -5) and (12, -5).

OpenStudy (anonymous):

@bronzegoddess helpppppp please)

OpenStudy (anonymous):

i'll try to :)

OpenStudy (anonymous):

@UnkleRhaukus , sorry james but i can't find my notes on this.. but the person I tagged is pretty good and this :)

OpenStudy (anonymous):

@bronzegoddess ok thanks anyways! and thanks for tagging him!

OpenStudy (anonymous):

welcome :)

OpenStudy (anonymous):

@bronzegoddess are u good at doing equations of graphs?

OpenStudy (anonymous):

yes, except hyperbolas and ellipces since its been some time..but i'll help you :)

OpenStudy (anonymous):

@4everjames you have a different question?

OpenStudy (anonymous):

@bronzegoddess lmao and im doing hyperbolas. and yes i have 8 questions all together that i need help with :(

OpenStudy (anonymous):

ish sorry, but i am at my grandma's and all my notes on hyperbolas are in another city :s i'll ask ash then (:

OpenStudy (anonymous):

@bronzegoddess ash?

OpenStudy (anonymous):

@ash2326

OpenStudy (anonymous):

@bronzegoddess ok cool.. thanks

OpenStudy (anonymous):

james, is there another name for co-vertices because i have done some research and people are just mention foci and vertix?

OpenStudy (anonymous):

@bronzegoddess The co-vertices are at the intersection of the minor axis and the ellipse.

OpenStudy (anonymous):

@bronzegoddess idk if there is another name for it

OpenStudy (anonymous):

okay, i'll do some more research then.

OpenStudy (anonymous):

@bronzegoddess i have an idea are u good at linear equations? id rather you help me with something you know how to do. i dont want you working to hard to help me

OpenStudy (anonymous):

yes, linear is pretty easy :)

OpenStudy (anonymous):

omg!!! @bronzegoddess ok we can do that instead. ill just skip parabolas for now.

OpenStudy (anonymous):

hyperbolas lol type away :)

OpenStudy (anonymous):

@bronzegoddess ahahaha oops! lol

OpenStudy (anonymous):

@bronzegoddess u ready?

OpenStudy (anonymous):

sure, shoot

OpenStudy (anonymous):

@bronzegoddess If a system has exactly one solution its called ?

OpenStudy (anonymous):

a monomial?

OpenStudy (anonymous):

@bronzegoddess consistent inconsistent independent dependent

OpenStudy (anonymous):

consistent

OpenStudy (anonymous):

@bronzegoddess thats what i put!!!! hahah *does dance* lol

OpenStudy (anonymous):

@bronzegoddess Solve the following system of equations by using the substitution method. 3y - 2x = 11 y + 2x = 9

OpenStudy (anonymous):

y=-2x+9, subsistute into the first then simplify

OpenStudy (anonymous):

@4everjames are you there?

OpenStudy (anonymous):

@bronzegoddess yes..

OpenStudy (anonymous):

can you simplify?

OpenStudy (anonymous):

@bronzegoddess (-2,-5)??

OpenStudy (anonymous):

you don't know how to simplify?

OpenStudy (anonymous):

@bronzegoddess yea.. just find a number that goes into both numerator and denominator right?

OpenStudy (anonymous):

3(-2x+9) -2x = 11 -6x +27-2x=11 27-11=8x x=2 then y=-2(2)+9=5 (2,5)

OpenStudy (anonymous):

i have to go

OpenStudy (anonymous):

@bronzegoddess will you be back?

OpenStudy (anonymous):

no, i have to study for an exam tmw... and its eggting late.. i'll help you another day sorry.. how many more question do you have?

OpenStudy (anonymous):

6 @bronzegoddess but its ok i appreciate the help that i did get! thanks so much